OpenStudy (anonymous):

$4 is what percent of $50?

OpenStudy (btaylor):

To determine percents, you use this formula: \[\frac{part}{whole} \times 100\] So:\[\frac{$4}{$50}\times 100 = 8percent\] So your answer is 8%